OpenStudy (anonymous):

PLEASE HELP, I DON'T UNDERSTAND THIS. MEDAL AND FAN FOR CORRECT ANSWER. Solve by multiplication and addition. Show each step of your work. 4x (minus) 3y = 11 3x + y = 18

OpenStudy (anonymous):

multiply 2nd eq with 3

OpenStudy (anonymous):

so its 4x - 3y =11 9x + 3y=54 so now add them toggether like (4x+9x)=65 the answer for x is 5 now substitute x in the first eq so its 4(5) - 3y = 11